4) Solve the system of equations using the substitution method
5x + 2y = 4 and 3x + y = 5



5) Solve the system of equations using either the substitution or the elimination method.
2x - 3y = -12 and - 4x + 6y = 15

Answer by stanbon(29603)

4) Solve the system of equations using the substitution method
5x + 2y = 4 and 3x + y = 5
Solve the 2nd for y: y=-3x+5
Substitute in the 1st: 5x+2(-3x+5)=4
5x-6x+10=4
-x=-6
x=6
--------
Substitute in y=-3x+5 to get:
y=-3(6)+5
y=-13
Solution: x=6, y=-13
-----------------
5) Solve the system of equations using either the substitution or the elimination method.
2x - 3y = -12 and - 4x + 6y = 15
Multiply the 1st by 2 to get:
4x-6y=-24
Add this to the 2nd to get:
0=-9
This is a contradiction and indicates there
is no solution for this system of equations.
Cheers,
Stan H.